Grade 12th passGeneral Physics

The radius of the nucleus of 8O16 is 3 × 10–15 m. Its density in kg/m³ will be about

Piyush Peeyush
10 Years agoGrade 12th pass

1 Answer

Suraj Prasad
10 Years ago
mass of 1 atom of O-16 is 16 a..m.u = 16*1.66*10^-27 Kg = 26.56 *10^-27 Kg.

Radius is given , so you can calculate the volume of it by doing V = 4/3 *Pi* r^3

Then we know density = mass/ Volume , apply it and get the answer.